1.FIVE TIMES A NUMBER IS 4 MORE THAN THE NUMBER. FIND THE NUMBER.
2.THE DIFFERENCE BETWEEN TWO NUMBERS IS 10 AND THEIR SUM IS 24. FIND THE NUMBERS

Answer by stanbon(48558)

1.FIVE TIMES A NUMBER IS 4 MORE THAN THE NUMBER. FIND THE NUMBER.
5x = x+4
4x=4
x=1
--------------

2.THE DIFFERENCE BETWEEN TWO NUMBERS IS 10 AND THEIR SUM IS 24. FIND THE NUMBERS
x-y=10
x+y=24
---------
Add the equations to get:
2x=34
x=17
Substitute to get:
17+y=24
y = 7
===========
Cheers,
Stan H.